what is the domain:
squareroot of 5-x /X^2+8X-9

Answer by stanbon(75887)

what is the domain:
squareroot of 5-x /X^2+8X-9
===
The denominator cannot be zero and
5-x cannot be negative.
---
Solve 5-x >= 0
x <=5
------
x^2+8x-9 = 0
(x+9)(x-1) = 0
x = -9 or x = 1
--------
On a number line graph x = -9, x=1 and the answer to x<=5
-----
So, x must be <=5 but not equal to 1 or -9
======================
Domain (-ing,-9)U(-9,1)U(1,5)
=============================
Cheers,
Stan H.
========